# Time-Scale Coupling Between States and Parameters in Recurrent Neural Networks

### v2

@@ Abstract
We study how gating mechanisms in recurrent neural networks (RNNs) implicitly induce adaptive learning-rate behavior, even when training is carried out with a fixed, global learning rate. This effect arises from the coupling between state-space time scales–parametrized by the gates–and parameter-space dynamics during gradient descent. By deriving exact Jacobians for leaky-integrator and gated RNNs, we obtain a first-order expansion that makes explicit how constant, scalar, and multi-dimensional gates reshape gradient propagation, modulate effective step sizes, and introduce anisotropy in parameter updates. These findings reveal that gates not only control information flow, but also act as data-driven preconditioners that adapt optimization trajectories in parameter space. We further draw formal analogies with learning-rate schedules, momentum, and adaptive methods such as Adam, pointing to possible redundancies. Empirical simulations corroborate these claims: in canonical synthetic sequence tasks (adding, copy) we show that gates induce lag-dependent effective learning rates and directional concentration of gradient flow, with multi-gate models matching or exceeding the anisotropic structure produced by Adam. These results highlight that optimizer-driven and gate-driven adaptivity are complementary but not equivalent mechanisms. Overall, this work provides a unified dynamical-systems perspective on how gating couples state evolution with parameter updates, explaining why gated architectures achieve robust trainability and stability in practice.
1 Introduction
Recurrent neural networks (RNNs) remain a central tool for modeling sequential data, yet their training dynamics continue to raise theoretical and practical challenges. A large body of work has shown how gating mechanisms—as in LSTMs and GRUs—help stabilize learning by regulating information flow in hidden states and mitigating gradient pathologies. From this perspective, gates are primarily understood as filters that control memory retention and short-term state evolution.
In this work, we demonstrate that gates have an additional, previously overlooked role: they also reshape the parameter dynamics during optimization. Specifically, we show that gating mechanisms implicitly induce adaptive learning-rate behavior, even when training is carried out with a fixed global step size. The same time-scale modulations that gates impose on state dynamics couple into parameter updates via backpropagation, altering effective step sizes, introducing lag dependence, and concentrating gradient flow into dominant directions. Thus, gates act not only as information filters but also as data-driven preconditioners of the optimization process.
Our theoretical analysis is based on a perturbative expansion of the Jacobian products arising in backpropagation. By deriving exact Jacobians for leaky-integrator, scalar-gated, and multi-gated RNNs, we make explicit how gates renormalize effective learning rates and introduce anisotropy in parameter updates. This perspective reveals structural analogies with optimization heuristics such as learning-rate schedules, momentum, and adaptive methods like Adam—yet these behaviors emerge endogenously from gating rather than being externally imposed.
We complement this analysis with controlled simulations that probe the predicted phenomena. In Section 7.1, we quantify how gates modulate lag-dependent effective learning rate profiles. In Section 7.2, we introduce an anisotropy index to measure how gates and optimizers shape the directional structure of gradient propagation. These experiments confirm that gate-induced corrections remain bounded but systematic, providing quantitative evidence that gates influence both the magnitude and orientation of temporal credit assignment. Section 7.3 discusses the broader implications of these findings for the trainability and stability of gated architectures.
Contributions.
We propose a unified dynamical-systems framework for analyzing how gating mechanisms affect the coupled dynamics of states and parameters in RNNs. Our contributions are:
@@ 1 Introduction
  • We show analytically that gates act as parametrized time scales that modulate lag-dependent effective learning rates.
@@
  • We extend the analysis to directional effects, introducing the anisotropy index to quantify how gates and optimizers shape the dominant subspaces of gradient propagation.
  • We validate these predictions through targeted simulations on canonical tasks (adding, copy), comparing leaky, scalar-gated, and multi-gated RNNs as well as adaptive optimization (Adam).
@@
Together, these results highlight that both gating and optimizer dynamics shape not only the magnitude but also the directional structure of temporal credit assignment, with task-dependent trade-offs.
Structure.
@@
The remainder of the paper is organized as follows. Section 2 surveys the most relevant literature on gradient pathologies, linear and state-space models, constrained recurrent dynamics, and gating mechanisms. Section 3 introduces the interpretation of RNN gating mechanisms as parametrized time scales and formalizes the specific models considered in our analysis. Section 4 reviews backpropagation through time in RNNs and revisits the vanishing/exploding gradient problem in this context. Section 5 presents our main theoretical contribution: the derivation of Jacobian matrices for the considered RNN models and the link between state-space time scales and parameter-space dynamics. Section 6 discusses well-known variants of gradient descent and suggests explicit connections between their behaviour and the effects induced by gating mechanisms. Section 7 complements the theory with simulations on well-known sequence tasks. Section 8 summarizes the key insights and outlines future research directions. The paper contains an Appendix with two sections. The first one, Appendix A describes the first-order expansion technique used in this paper and provides numerical examples justifying its robustness. The second one, Appendix B provides a description of the well-know Adam algorithm.
2 Related Works
@@
The study of recurrent neural networks (RNNs) has long been shaped by the well-known vanishing and exploding gradient problem [24], which affects the stability and learnability of long-term dependencies. While this issue traditionally focuses on the magnitude of gradients, recent work by Zucchet and Orvieto [38] highlights a complementary phenomenon: as the memorization capacity of an RNN increases during training, the network output can become highly sensitive to small parameter variations. This heightened sensitivity arises even in the absence of vanishing or exploding gradients, suggesting that RNNs may undergo abrupt changes in behavior due to intrinsic instabilities in the learned parameter–state mapping. In a related vein, Ceni [3] proposed Random Orthogonal Additive Filters, a class of architectures designed to stabilize and enrich recurrent dynamics via orthogonal transformations, mitigating instability without sacrificing representational power.
@@
Several lines of research have explored simplified linear models as a means of studying and designing recurrent dynamics with greater interpretability. Continuous-time state-space models, such as those in Gu et al. [9, 10], have enabled efficient training of sequence models with long-range dependencies by leveraging structured state-space kernels. More recently, Muca Cirone et al. [23] provided a theoretical analysis of linear state-space models (LSSMs), clarifying their expressive power and the role of parameterization in controlling stability and memory retention. These models offer a valuable bridge between classical control theory and modern deep learning approaches to sequence modeling.
@@
Gating mechanisms have received particular attention for their role in modulating information flow and improving trainability. From a theoretical standpoint, Chen et al. [5] and Gilboa et al. [8] analyzed the effects of gating on dynamical isometry and mean-field properties in RNNs, showing that gates can help preserve gradient flow and condition the optimization landscape. Related analyses in feedforward networks [25] underline the generality of dynamical isometry principles across architectures. Architecturally, stacked gated RNNs [32] and hybrid designs such as the “Just Another Network” (JANET) [33] demonstrate the versatility of gating in achieving both memory retention and efficient optimization. More recent theoretical work further strengthens this connection: Krishnamurthy et al. [20] and Can et al. [2] showed that gates can create slow modes in recurrent dynamics, directly modulating Jacobian spectra and shaping effective memory timescales. Empirical studies such as Quax et al. [26] confirmed that RNNs tend to develop adaptive timescales when trained on multi-scale sequential data, further illustrating the deep connection between gating and temporal structure.
In contrast to these prior works, which often study either stability, memory, or optimization in isolation, our work examines the joint dynamics of states and parameters under the influence of gating. By framing gating as a parametrized time-scale modulation and applying perturbative analysis to the Jacobian products that arise in backpropagation, we connect architectural design choices directly to their effect on parameter update anisotropy, gradient scaling, and memory retention across time scales.
3 RNNs and Time-Scales
@@
3.1 From continuous to discrete time
@@
$x_{t+1}=\phi\!\left(W^{r}x_{t}+W^{i}u_{t}\right),$ (5)
with discrete-time output
$z_{t}=\psi(x_{t}),$ (6)
and, in the linear case,
$z_{t}=W^{o}x_{t}.$ (7)
3.2 Global time rescaling
@@
3.3 General time warping and gating
@@
$=g(t)\,\phi\!\left(W^{r}x(t)+W^{i}u(t)\right)-g(t)\,x(t),$ (12)
where
$g(t):=\frac{dc(t)}{dt}$ (13)
is the instantaneous state-update rate. When $g(t)\equiv\alpha$, we recover the leaky case (9).
@@
where $\sigma(\cdot)\in(0,1)$ is the logistic sigmoid, $W^{r,g}\in\mathbb{R}^{1\times N_{r}}$, and $W^{i,g}\in\mathbb{R}^{1\times N_{i}}$ are gate parameters. This parametrization ensures that the update rate $g(t)$ remains bounded in $(0,1)$, thereby smoothly interpolating between two extremes: $g(t)\approx 0$ corresponds to almost frozen dynamics (long time scale), while $g(t)\approx 1$ yields rapid updates (short time scale). Because $g(t)$ depends both on the current state $x(t)$ and the input $u(t)$, the effective time scale of the network becomes data- and state-dependent. This means that the model can adaptively modulate its temporal resolution on the fly, allocating fast or slow dynamics depending on the context, rather than relying on a fixed global constant $\alpha$ as in the leaky-integrator case.
Discretizing (12) with $\delta t=1$ gives
@@
which makes the dynamics invariant to a global rescaling of time and, more generally, allows them to adapt to smoothly varying changes in the time axis through the gate $g_{t}$.
@@
where $\odot$ denotes elementwise multiplication and $g_{t}\in(0,1)^{N_{r}}$. Here $\sigma(\cdot)$ in (14) is applied componentwise, with $W^{r,g}\in\mathbb{R}^{N_{r}\times N_{r}}$ and $W^{i,g}\in\mathbb{R}^{N_{r}\times N_{i}}$. Each neuron thus possesses its own time-dependent update rate (or time scale), enabling fine-grained adaptation of the network’s temporal dynamics.
4 BPTT-based training of RNNs
4.1 Stochastic gradient descent in neural networks
@@
where $l$ indexes the training iteration, $\mu>0$ is the learning rate (or step size), and $\mathcal{E}_{t}$ is the loss at sample $t$ in the mini-batch. The stochasticity arises because each mini-batch is drawn at random from the training set, leading to noisy gradient estimates. For instance, in a regression setting one may use the squared error
$\mathcal{E}_{t}\;=\;\lVert y_{t}-z_{t}\rVert^{2},$
where $y_{t}$ is the target output, $z_{t}$ is the model prediction, and $\theta$ collects all trainable parameters of the network.
@@
Because the error gradients must be propagated backward through the sequence of states, this procedure is known as backpropagation. When applied to recurrent neural networks, where parameters are shared across time steps, the method is referred to as backpropagation through time (BPTT) [36].
4.2 Vanishing and exploding gradients
@@
$\frac{\partial x_{t}}{\partial x_{k}}$ $=\prod_{j=k+1}^{t}J_{j},$ (19)
$J_{j}$ $=\frac{\partial x_{j}}{\partial x_{j-1}},$ (20)
@@
Using the submultiplicative property of matrix norms,
$\left\lVert\frac{\partial x_{t}}{\partial x_{k}}\right\rVert\;\leq\;\prod_{j=k+1}^{t}\left\lVert J_{j}\right\rVert,$ (21)
and substituting into (18) gives the following bound on the gradient norm:
@@
Consider now the standard discrete-time RNN model (5). The Jacobian at step $j$ has the form
$J_{j}=D_{j-1}W^{r},$ (23)
where
$D_{j}=\mathrm{diag}(\phi^{\prime}(a_{j}))$ (24)
is a diagonal matrix containing the derivatives of the activation function $\phi$ evaluated at the pre-activation
$a_{j-1}=W^{r}x_{j-1}+W^{i}u_{j-1}.$ (25)
In this case, (21) specializes to
$\prod_{j=k+1}^{t}\left\lVert D_{j-1}\right\rVert\,\left\lVert W^{r}\right\rVert.$ (26)
From (26), it is clear that if the spectral norm $\lVert W^{r}\rVert$ or the activation derivatives $\lVert D_{j-1}\rVert$ are consistently smaller than $1$, the product will decay exponentially with (tk)(t-k), leading to vanishing gradients. Conversely, if either factor is consistently greater than $1$, the product will grow exponentially, producing exploding gradients. This spectral-norm dependence of the gradient magnitude is well documented in the literature [24] and motivates many of the stability constraints and reparameterizations used in modern RNN architectures.
5 How time-scales in RNN states affect parameter dynamics
In Section 5.1, we derive the exact expressions for the Jacobian matrices associated with the RNN variants introduced in Section 3. These Jacobians govern how perturbations to the hidden state propagate through time, and thus play a central role in both forward signal evolution and backward gradient flow. Building on these results, Section 5.2 analyzes how the time-scales embedded in the state-space dynamics interact with the optimization process. In particular, we show how the presence of constant, scalar, or multiple gating mechanisms influences the effective learning rate and the direction of parameter updates, thereby shaping the behaviour of the gradient descent algorithm.
5.1 Jacobian matrices
5.1.1 Leaky-integrator neurons
@@
where $I$ is the identity matrix and $D_{j-1}=\mathrm{diag}(\phi^{\prime}(a_{j-1}))$ (see (24)). Relative to the standard RNN Jacobian $D_{j-1}W^{r}$, the constant gate $\alpha$ scales the recurrent contribution while adding a skip connection proportional to $(1-\alpha)I$, thereby tempering gradient decay or explosion.
5.1.2 Single scalar gate
Consider now the scalar-gated model (15). Defining the gate pre-activation
$a^{g}_{t}=W^{r,g}x_{t}+W^{i,g}u_{t},\ a^{g}_{t}\in\mathbb{R},$ (28)
the Jacobian becomes
$J_{j}$ $=\Big{[}\phi(a_{j-1})\frac{\partial g_{j-1}}{\partial x_{j-1}}+g_{j-1}D_{j-1}W^{r}\Big{]}+\Big{[}x_{j-1}\frac{\partial(1-g_{j-1})}{\partial x_{j-1}}+(1-g_{j-1})I\Big{]}$ (29)
$=\big{(}\phi(a_{j-1})-x_{j-1}\big{)}J^{g}_{j-1}+g_{j-1}D_{j-1}W^{r}+(1-g_{j-1})I$ (30)
$=G_{j-1}+g_{j-1}D_{j-1}W^{r}+(1-g_{j-1})I,$ (31)

### v3

@@ Abstract
We study how gating mechanisms in recurrent neural networks (RNNs) implicitly induce adaptive learning-rate behavior, even when training is carried out with a fixed, global learning rate. This effect arises from the coupling between state-space time scales–parametrized by the gates–and parameter-space dynamics during gradient descent. By deriving exact Jacobians for leaky-integrator and gated RNNs, we obtain a first-order expansion that makes explicit how constant, scalar, and multi-dimensional gates reshape gradient propagation, modulate effective step sizes, and introduce anisotropy in parameter updates. These findings reveal that gates not only control information flow, but also act as data-driven preconditioners that adapt optimization trajectories in parameter space. We further draw formal analogies with learning-rate schedules, momentum, and adaptive methods such as Adam. Empirical simulations corroborate these claims: in several sequence tasks, we show that gates induce lag-dependent effective learning rates and directional concentration of gradient flow, with multi-gate models matching or exceeding the anisotropic structure produced by Adam. These results highlight that optimizer-driven and gate-driven adaptivity are complementary but not equivalent mechanisms. Overall, this work provides a unified dynamical systems perspective on how gating couples state evolution with parameter updates, explaining why gated architectures achieve robust trainability and stability in practice.
Index Terms:
Recurrent neural networks; Dynamical systems; Gates; Time scale coupling; Gradient descent.
I Introduction
Training recurrent neural networks (RNNs) is typically analyzed from two largely separate perspectives. On the one hand, state-space dynamics emphasize how gating mechanisms stabilize hidden trajectories, regulate memory retention, and mitigate vanishing or exploding gradients. On the other hand, parameter-space dynamics emphasize the role of optimization algorithms—such as momentum or Adam—in adapting learning rates and reshaping update directions to ease training. These perspectives are usually studied in isolation: gates are viewed as filters of information flow, optimizers as external controllers of learning dynamics.
What has been missing is a unified account of how these two domains interact. In particular, why do gated RNNs often train stably? This robustness suggests that gates must influence not only the flow of information in state space, but also the dynamics of parameter updates. Yet the mechanisms by which state-time scales couple into optimization have remained implicit.
In this work, we show that gating mechanisms also act as structural priors for optimization. By parametrizing time scales in state dynamics, gates reshape the Jacobian products that govern backpropagation, thereby inducing lag-dependent effective learning rates and directional anisotropy in parameter updates. From this perspective, gates are not just stabilizers of state trajectories: they also function as implicit, data-driven preconditioners of the optimization process. This reveals formal analogies with well-known optimization heuristics such as learning-rate schedules, momentum, and Adam—but here these behaviors emerge endogenously from gating rather than being externally imposed.
Our contributions are:
@@ I Introduction
  • We show analytically that gates act as parametrized time scales that modulate lag-dependent effective learning rates. Our analysis is based on a perturbative expansion of Jacobian products, which makes explicit how constant, scalar, and multi-gate RNNs modulate effective learning rates.
@@
  • We extend the theoretical analysis to directional effects, introducing the anisotropy index to quantify how gates shape the dominant subspaces of gradient propagation.
@@
  • We validate these predictions through targeted simulations on canonical tasks. Our results highlight that both gating and optimizer dynamics shape not only the magnitude but also the directional structure of temporal credit assignment, with task-dependent trade-offs.
@@
The remainder of the paper is organized as follows. Section II surveys the most relevant literature. Section III introduces the interpretation of RNN gating mechanisms as parametrized time scales and formalizes the specific models considered in our analysis. Section IV reviews backpropagation through time in RNNs and revisits the vanishing/exploding gradient problem in this context. Section V presents our main theoretical contribution: the derivation of Jacobian matrices for the considered RNN models and the link between state-space time scales and parameter-space dynamics. Section VI discusses well-known variants of gradient descent and suggests explicit connections between their behaviour and the effects induced by gating mechanisms. Section VII complements the theory with simulations on well-known sequence tasks. Section VIII summarizes the key insights and outlines future research directions. The paper contains Supplementary Material, including details about the first-order expansion technique used in the paper, a summary of the well-known Adam algorithm, and all figures related to the simulations.
II Related Works
@@
The study of RNNs has long been shaped by the well-known vanishing and exploding gradient problem [1], which affects the stability and learnability of long-term dependencies. While this issue traditionally focuses on the magnitude of gradients, recent work by Zucchet et al. [2] highlights a complementary phenomenon: as the memorization capacity of an RNN increases during training, the network output can become highly sensitive to small parameter variations. This heightened sensitivity arises even in the absence of vanishing or exploding gradients, suggesting that RNNs may undergo abrupt changes in behavior due to intrinsic instabilities in the learned parameter–state mapping. In a related vein, Ceni [3] proposed Random Orthogonal Additive Filters, a class of architectures designed to stabilize and enrich recurrent dynamics via orthogonal transformations, mitigating instability without sacrificing representational power.
@@
Several lines of research have explored simplified linear models as a means of studying and designing recurrent dynamics with greater interpretability. Continuous-time state-space models, such as those in [4, 5], have enabled efficient training of sequence models with long-range dependencies by leveraging structured state-space kernels. More recently, Muca et al. [6] provided a theoretical analysis of linear state-space models, clarifying their expressive power and the role of parameterization in controlling stability and memory retention. These models offer a valuable bridge between classical control theory and modern deep learning approaches to sequence modeling.
@@
Gating mechanisms have received particular attention for their role in modulating information flow and improving trainability. From a theoretical standpoint, [21] and [22] analyzed the effects of gating on dynamical isometry and mean-field properties in RNNs, showing that gates can help preserve gradient flow and condition the optimization landscape. Related analyses in feedforward networks [23] underline the generality of dynamical isometry principles across architectures. Architecturally, stacked gated RNNs [24] and hybrid designs such as the “Just Another Network” [25] demonstrate the versatility of gating in achieving both memory retention and efficient optimization. More recent theoretical work further strengthens this connection: Krishnamurthy et al. [26] and Can et al. [27] showed that gates can create slow modes in recurrent dynamics, directly modulating Jacobian spectra and shaping effective memory timescales. Empirical studies such as [28] confirmed that RNNs tend to develop adaptive timescales when trained on multi-scale sequential data, further illustrating the deep connection between gating and temporal structure.
III RNNs and Time-Scales
@@
III-A From continuous to discrete time
@@
$x_{t+1}=\phi\!\left(W^{r}x_{t}+W^{i}u_{t}\right).$ (5)
III-B Global time rescaling
@@
III-C General time warping and gating
@@
$=g(t)\,\phi\!\left(W^{r}x(t)+W^{i}u(t)\right)-g(t)\,x(t),$
where $g(t):=\frac{dc(t)}{dt}$ is the instantaneous state-update rate at time $t$. When $g(t)\equiv\alpha$, we recover the leaky case (7).
@@
where $\sigma(\cdot)\in(0,1)$ is the logistic sigmoid, $W^{r,g}\in\mathbb{R}^{1\times N_{r}}$, and $W^{i,g}\in\mathbb{R}^{1\times N_{i}}$ are gate parameters. This parametrization ensures that the update rate $g(t)$ remains bounded in $(0,1)$. Because $g(t)$ depends both on the current state $x(t)$ and the input $u(t)$, the effective time scale of the network becomes data- and state-dependent.
Discretizing (LABEL:eq:rnn_ct_warping) with $\delta t=1$ gives
@@
which makes the dynamics invariant to a global rescaling of time via $g_{t}$.
@@
where $\odot$ denotes elementwise multiplication and $g_{t}\in(0,1)^{N_{r}}$. Here $\sigma(\cdot)$ in (10) is applied componentwise, with $W^{r,g}\in\mathbb{R}^{N_{r}\times N_{r}}$ and $W^{i,g}\in\mathbb{R}^{N_{r}\times N_{i}}$. Each neuron thus possesses its own update rate (or time scale), enabling fine-grained adaptation of the network’s temporal dynamics.
IV BPTT-based training of RNNs
IV-A Stochastic gradient descent in neural networks
@@
where $l$ indexes the training iteration, $\mu>0$ is the learning rate (or step size), and $\mathcal{E}_{t}$ is the loss at sample $t$ in the mini-batch. The stochasticity arises because each mini-batch is drawn at random from the training set, leading to noisy gradient estimates.
@@
Because the gradients must be propagated backward through the sequence of states, this procedure is known as backpropagation. When applied to recurrent neural networks, where parameters are shared across time steps, the method is referred to as backpropagation through time (BPTT) [31].
IV-B Vanishing and exploding gradients
@@
$\frac{\partial x_{t}}{\partial x_{k}}=\prod_{j=k+1}^{t}J_{j},\ J_{j}=\frac{\partial x_{j}}{\partial x_{j-1}},$ (15)
@@
Using the submultiplicative property of matrix norms, $\left\lVert\frac{\partial x_{t}}{\partial x_{k}}\right\rVert\;\leq\;\prod_{j=k+1}^{t}\left\lVert J_{j}\right\rVert,$ and substituting into (14) gives the following upper bound on the gradient norm:
@@
Consider now the standard discrete-time RNN model (5). The Jacobian at step $j$ has the form $J_{j}=D_{j-1}W^{r}$, where $D_{j}=\mathrm{diag}(\phi^{\prime}(a_{j}))$ is a diagonal matrix containing the derivatives of the activation function $\phi$ evaluated at the pre-activation $a_{j-1}=W^{r}x_{j-1}+W^{i}u_{j-1}$. The upper bound specializes to $\prod_{j=k+1}^{t}\left\lVert D_{j-1}\right\rVert\,\left\lVert W^{r}\right\rVert$. It is clear that if the spectral norm $\lVert W^{r}\rVert$ or the activation derivatives $\lVert D_{j-1}\rVert$ are consistently smaller than $1$, the product will decay exponentially with (tk)(t-k), leading to vanishing gradients. Conversely, if either factor is consistently greater than $1$, the product will grow exponentially, producing exploding gradients. This spectral-norm dependence of the gradient magnitude is well documented in the literature [1] and motivates many of the stability constraints and reparameterizations used in modern RNN architectures.
V How time-scales in RNN states affect parameter dynamics
In Section V-A, we derive the exact expressions for the Jacobian matrices associated with the RNN variants introduced in Section III. These Jacobians govern how perturbations to the hidden state propagate through time, and thus play a central role in both forward signal evolution and backward gradient flow. Building on these results, Section V-B analyzes how the time-scales embedded in the state-space dynamics interact with the optimization process. In particular, we show how the presence of constant, scalar, or multiple gating mechanisms influences the effective learning rate and the direction of parameter updates, thereby shaping the behaviour of the gradient descent algorithm.
V-A Jacobian matrices
V-A1 Leaky-integrator neurons
@@
where $I$ is the identity matrix and $D_{j-1}=\mathrm{diag}(\phi^{\prime}(a_{j-1}))$. Relative to the standard RNN Jacobian $D_{j-1}W^{r}$, the constant gate $\alpha$ scales the recurrent contribution while adding a skip connection proportional to $(1-\alpha)I$, thereby tempering gradient decay or explosion.
V-A2 Single scalar gate
Consider now the scalar-gated model (11). Defining the gate pre-activation $a^{g}_{t}=W^{r,g}x_{t}+W^{i,g}u_{t},\ a^{g}_{t}\in\mathbb{R}$, the Jacobian becomes
$J_{j}$ $=\Big{[}\phi(a_{j-1})\frac{\partial g_{j-1}}{\partial x_{j-1}}+g_{j-1}D_{j-1}W^{r}\Big{]}$ (18)
$+\Big{[}x_{j-1}\frac{\partial(1-g_{j-1})}{\partial x_{j-1}}+(1-g_{j-1})I\Big{]}$
$=\big{(}\phi(a_{j-1})-x_{j-1}\big{)}J^{g}_{j-1}+g_{j-1}D_{j-1}W^{r}+(1-g_{j-1})I$
$=G_{j-1}+g_{j-1}D_{j-1}W^{r}+(1-g_{j-1})I,$